Tom is walking 12 miles from A to B.
He already walked x miles.
He needs to walk y miles more.

Found 3 solutions by Boreal, ikleyn, ewatrrr:
Answer by Boreal(15235)

Not clear what you want.
x+y=12
he has 12-x miles to walk
y=12-x with units miles

Answer by ikleyn(52803)

.

This message is exclusively interesting, but . . .


But without your question, it does not make sense.


You may post this exciting story to New York Times as an essay.


I do not recommend you post the "problems" to the forum with no question,
because it may happen that I will delete or destroy such a message.



At this forum,  we do not accept the posts with no question.

The problems with no questions,  address to this web-site

www.mad-house_Homework_help.com

They accept such posts for free  (and with no objections).



Answer by ewatrrr(24785)


Hi
12mi from A to B
walked 'x' miles
remaining is 'y' miles
 y = 12-x
and 
x + 12-x = 12
For ex: IF x = 5mi ⇒ y = 12-x = 7
Using the variable like x and y helps one solve problems!
Wish You the Best in your Studies.